# Complexity of deep computations via topology of function spaces

### v1

@@ 4.1. A measure-theoretic version of NIP
Given a Hausdorff space $X$ and a measurable space $(Y,\Sigma)$, we say that $f:X\rightarrow Y$ is universally measurable (with respect to $\Sigma$) if $f^{-1}(E)$ is Borel for every $E\in\Sigma$, i.e., $f^{-1}(E)$ is $\mu$-measurable for every Radon measure $\mu$ on $X$. When $Y=\mathbb{R}$ we will always take $\Sigma=\mathcal{B}(\mathbb{R})$, the Borel $\sigma$-algebra of $\mathbb{R}$.
@@
Intuitively, a function is universally measurable if it is “measurable no matter which reasonable way you try to measure things on its domain”. The concept of universal measurability emerged from work of Kallianpur and Sazonov, in the late 1950’s and 1960s, with later developments by Blackwell, Darst, and others, building on earlier ideas of Gnedenko and Kolmogorov from the 1950s. See [Pap:2002, Chapters 1 and 2].
@@
Following [BFT_1978_PCompactBaire], the collection of all universally measurable real-valued functions will be denoted by $M_{r}(X)$.
In the context of deep computations, we will be interested in transition maps of a state space $L\subseteq\mathbb{R}^{\mathcal{P}}$ into itself. There are two natural $\sigma$-algebras one can consider in the product space $\mathbb{R}^{\mathcal{P}}$: the Borel $\sigma$-algebra, i.e., the $\sigma$-algebra generated by open sets in $\mathbb{R}^{\mathcal{P}}$, and the cylinder $\sigma$-algebra, i.e., the $\sigma$-algebra generated by the sub-basic open sets in $\mathbb{R}^{\mathcal{P}}$. Note that when $\mathcal{P}$ is countable, both $\sigma$-algebras coincide, but in general the cylinder $\sigma$-algebra is strictly smaller. We will use the cylinder $\sigma$-algebra to define universally measurable maps $f:\mathbb{R}^{\mathcal{P}}\rightarrow\mathbb{R}^{\mathcal{P}}$. The reason for this choice is the following characterization:
Lemma 4.4.
Let $X$ be a Hausdorff space and $Y=\prod_{i\in I}Y_{i}$ be any product of measurable spaces $(Y_{i},\Sigma_{i})$ for $i\in I$. Let $\Sigma_{Y}$ be the cylinder $\sigma$-algebra generated by the measurable spaces $(Y_{i},\Sigma_{i})$. Let $f:X\rightarrow Y$. The following are equivalent:
@@
The preceding lemma says that a transition map is universally measurable if and only if it is universally measurable on all its features; in other words, we can check measurability of a transition just by checking measurability feature by feature. We will denote by $M_{r}(X,\mathbb{R}^{\mathcal{P}})$ the collection of all universally measurable functions $f:X\rightarrow\mathbb{R}^{\mathcal{P}}$ (with respect to the cylinder $\sigma$-algebra), endowed with the topology of pointwise convergence.
We will need the following result about NIP and universally measurable functions:
Theorem 4.5 (Bourgain-Fremlin-Talagrand, Theorem 2F in [BFT_1978_PCompactBaire]).
@@
4.2. Universal Monte Carlo computability of deep computations
We now wish to define the concept of a deep computation being computable except on a set of arbitrarily small measure “no matter which reasonable way you try to measure things on its domain” (see the remarks following definition 4.1). This is the concept of universal Monte Carlo computability defined below (Definition 4.6). To motivate the definition, we need to recall two facts:
@@ 4.2. Universal Monte Carlo computability of deep computations
    Littlewoood’s second principle states that every Lebesgue measurable function is “nearly continuous”. The formal version of this, which is Luzin’s theorem, states that if $(X,\Sigma,\mu)$ a Radon measure space and $Y$ be a second-countable topological space (e.g., $Y=\mathbb{R}^{\mathcal{P}}$ with $\mathcal{P}$ countable ) equipped with a Borel algebra, then any given $f:X\to Y$ is measurable if and only if for every $E\in\Sigma$ and every $\varepsilon>0$ there exists a closed $F\subseteq E$ such that the restriction $f|F$ is continuous.
@@
    Computability of deep computations can is characterized in terms of continuous extendibility of computations. This is at the core of [alva2024approximability].
@@
Let $(L,\mathcal{P},\Gamma)$ be a CCS. We say that a transition $f:L\rightarrow L$ is universally Monte Carlo computable if and only if there exists $\tilde{f}:\mathcal{L}_{\text{sh}}\rightarrow\mathcal{L}_{\text{sh}}$ extending $f$ such that for every sizer $r_{\bullet}$ there is a sizer $s_{\bullet}$ such that the restriction $\tilde{f}|_{\mathcal{L}[r_{\bullet}]}:\mathcal{L}[r_{\bullet}]\rightarrow\mathcal{L}[s_{\bullet}]$ is universally measurable, i.e., $\pi_{P}\circ\tilde{f}|_{\mathcal{L}[r_{\bullet}]}:\mathcal{L}[r_{\bullet}]\rightarrow[-s_{P},s_{P}]$ is $\mu$-measurable for every Radon probability measure $\mu$ on $\mathcal{L}[r_{\bullet}]$.
4.3. Bourgain-Fremlin-Talagrand, NIP, and universal Monte Carlo computability of deep computations
Theorem 4.5 immediately yields the following.
@@ 4.3. Bourgain-Fremlin-Talagrand, NIP, and universal Monte Carlo computability of deep computations
Let $(L,\mathcal{P},\Gamma)$ be a CCS satisfying the Extendability Axiom. Let $R$ be an exhaustive collection of sizers. Let $\Delta\subseteq\Gamma$ be $R$-confined. If $\pi_{P}\circ\Delta|_{L[r_{\bullet}]}$ satisfies the NIP for all $P\in\mathcal{P}$ and all $r_{\bullet}\in R$, then every deep computation is universally Monte Carlo computable.
@@
By the Extendability Axiom, Theorem 4.5 and Lemma 1.13 we have $\overline{\pi_{P}\circ\tilde{\Delta}|_{\mathcal{L}[r_{\bullet}]}}\subseteq M_{r}(\mathcal{L}[r_{\bullet}])$ for all $r_{\bullet}\in R$ and $P\in\mathcal{P}$. Let $f\in\overline{\Delta}$ be a deep computation. Write $f=\mathcal{U}\lim_{i}\gamma_{i}$ as an ultralimit of computations in $\Delta$. Define $\tilde{f}:=\mathcal{U}\lim_{i}\tilde{\gamma}_{i}$. Then, for all $r_{\bullet}\in R$ and $P\in\mathcal{P}$ $\pi_{P}\circ\tilde{\gamma}_{i}|_{\mathcal{L}[r_{\bullet}]}\in M_{r}(\mathcal{L}[r_{\bullet}])$ for all $i$, so $\pi_{P}\circ f|_{\mathcal{L}[r_{\bullet}]}\in\overline{\pi_{P}\circ\tilde{\Delta}|_{\mathcal{L}[r_{\bullet}]}}\subseteq M_{r}(\mathcal{L}[r_{\bullet}])$. ∎
Question 4.8.
@@
4.4. Talagrand stability, Fremlin’s dichotomy, NIP, and universal Monte Carlo computability of deep computations
There is another notion closely related to NIP, introduced by Talagrand in [talagrand1984pettis] while studying Pettis integration. Suppose that $X$ is a compact Hausdorff space and $A\subseteq\mathbb{R}^{X}$. Let $\mu$ be a Radon probability measure on $X$. Given a $\mu$-measurable set $E\subseteq X$, a positive integer $k$ and real numbers $a<b$. we write:
$D_{k}(A,E,a,b)=\bigcup_{f\in A}\{x\in E^{2k}:f(x_{2i})\leq a,\hskip 2.84526ptf(x_{2i+1})\geq b\hskip 2.84526pt{\text{ for all }i<k}\}$
@@
where $\mu*$ denotes the outer measure (we work with outer since the sets $D_{k}(A,E,a,b)$ need not be $\mu$-measurable). This is certainly the case when $A$ is a countable set of continuous (or $\mu$-measurable) functions.
Notation 4.9.
For a measure $\mu$ on a set $X$, the set of all $\mu$-measurable functions will denoted by $\mathscr{M}^{0}(X,\mu)$.
The following lemma establishes that Talagrand stability is a way to ensure that deep computations are definable by measurable functions.
@@
First, observe that a subset of a $\mu$-stable set is $\mu$-stable. To show that $\overline{A}$ is $\mu$-stable, observe that $D_{k}(\overline{A},E,a,b)\subseteq D_{k}(A,E,a^{\prime},b^{\prime})$ where $a<a^{\prime}<b^{\prime}<b$ and $E$ is a $\mu$-measurable set with positive measure. It suffices to show that $\overline{A}\subseteq\mathscr{M}^{0}(X,\mu)$. Suppose that there exists $f\in\overline{A}$ such that $f\notin\mathscr{M}^{0}(X,\mu)$. By a characterization of measurable functions (see 413G in [fremlin2003vol4]), there exists a $\mu$-measurable set $E$ of positive measure and $a<b$ such that $\mu^{*}(P)=\mu^{*}(Q)=\mu(E)$ where $P=\{x\in E:f(x)\leq a\}$ and $Q=\{x\in E:f(x)\geq b\}$. Then, for any $k\geq 1$: $(P\times Q)^{k}\subseteq D_{k}(\{f\},E,a,b)$ so $(\mu^{2k})^{*}(D_{k}(\{f\},E,a,b))=(\mu^{*}(P)\mu^{*}(Q))^{k}=(\mu(E))^{2k}$. Thus, $\{f\}$ is not $\mu$-stable, but we argued before that a subset of a $\mu$-stable set must be $\mu$-stable. ∎
We say that $A$ is universally Talagrand stable if $A$ is Talagrand $\mu$-stable for every Radon probability measure $\mu$ on $X$. An argument similar to the proof of 4.5, yields the following:
Theorem 4.11.
Let $(L,\mathcal{P},\Gamma)$ be a CCS satisfying the Extendability Axiom. If $\pi_{P}\circ\Delta|_{L[r_{\bullet}]}$ is universally Talagrand stable for all $P\in\mathcal{P}$ and all sizers $r_{\bullet}$, then every deep computation is universally Monte Carlo computable.
It is then natural to ask: what is the relationship between Talagrand stability and the NIP? The following dichotomy will be useful.
Lemma 4.12 (Fremlin’s Dichotomy, 463K in [fremlin2003vol4]).
If $(X,\Sigma,\mu)$ is a perfect $\sigma$-finite measure space (in particular, for $X$ compact and $\mu$ a Radon probability measure on $X$) and $\{f_{n}:n\in\mathbb{N}\}$ be a sequence of real-valued measurable functions on $X$, then one (and only one) of the following conditions holds:
  1. (i)
    $\{f_{n}:n\in\mathbb{N}\}$ has a subsequence that converges $\mu$-almost everywhere,
  2. (ii)
    $\{f_{n}:n\in\mathbb{N}\}$ has a subsequence with no $\mu$-measurable accumulation point in $\mathbb{R}^{X}$.
The preceding lemma can be considered as a measure-theoretic version of Rosenthal’s dichotomy. Combining this dichotomy with Theorem 4.5, we get the following result:
Theorem 4.13.
  1. (i)
    $\overline{A}\subseteq M_{r}(X)$.
  2. (ii)
    For every compact $K\subseteq X$, $A|_{K}$ satisfies the NIP.
  3. (iii)
  4. (iv)
    For every Radon measure $\mu$ on $X$ and every sequence $\{f_{n}:n\in\mathbb{N}\}\subseteq A$, there is a subsequence that converges $\mu$-almost everywhere.
Proof.
Notice that the equivalence (i)-(iii) is Theorem 4.5. Notice that the equivalence of (iii) and (iv) is Fremlin’s Dichotomy (Theorem 4.12). ∎
Finally, it is natural to ask what the connection is between Talagrand stability and NIP.
@@
By Theorem 4.5, it suffices to show that $A$ is relatively countably compact in $\mathscr{M}^{0}(X,\mu)$ for all Radon probability measure $\mu$ on $X$. Since $A$ is Talagrand $\mu$-stable for any such $\mu$, we have $\overline{A}\subseteq\mathscr{M}^{0}(X,\mu)$. In particular, $A$ is relatively countably compact in $\mathscr{M}^{0}(X,\mu)$. ∎
Question 4.15.
Is the converse true?
The following two results suggest that the precise connection between Talagrand stability and NIP may be sensitive to set-theoretic axioms (even assuming countability of $A$).
Theorem 4.16 (Talagrand, Theorem 9-3-1(a) in [talagrand1984pettis]).
@@
Theorem 4.17 (Fremlin, Shelah, [fremlin1993pointwise]).

### v2

@@ 4.1. NIP and Monte Carlo computability of deep computations
Given a Hausdorff space $X$ and a measurable space $(Y,\Sigma)$, we say that $f:X\rightarrow Y$ is universally measurable (with respect to $\Sigma$) if $f^{-1}(E)$ is $\mu$-measurable for every Radon measure $\mu$ on $X$ and every $E\in\Sigma$. When $Y=\mathbb{R}$, we will always take $\Sigma=\mathcal{B}(\mathbb{R})$, the Borel $\sigma$-algebra of $\mathbb{R}$.
@@
Intuitively, a function is universally measurable if it is “measurable no matter which reasonable way you try to measure things on its domain”. The concept of universal measurability emerged from work of Kallianpur and Sazonov, in the late 1950’s and 1960s — with later developments by Blackwell, Darst and others — building on earlier ideas of Gnedenko and Kolmogorov from the 1950s. See [PAP02, Chapters 1 and 2].
@@
Following [BFT78], the collection of all universally measurable real-valued functions on $X$ will be denoted by $M_{r}(X)$. Given a fixed Radon measure $\mu$ on $X$, the collection of all $\mu$-measurable real-valued functions on $X$ will be denoted by $\mathscr{M}^{0}(X,\mu)$.
In the context of deep computations, we are interested in transition maps of a state space $L\subseteq\mathbb{R}^{\mathcal{P}}$ into itself. There are two natural $\sigma$-algebras one can consider in the product space $\mathbb{R}^{\mathcal{P}}$: the Borel $\sigma$-algebra, i.e., the $\sigma$-algebra generated by open sets in $\mathbb{R}^{\mathcal{P}}$, and the cylinder $\sigma$-algebra (i.e., the $\sigma$-algebra generated by the sub-basic open sets in $\mathbb{R}^{\mathcal{P}}$, that is, the set $\pi_{i}^{-1}(U)$ with $U\subseteq\mathbb{R}$ open). Note that when $\mathcal{P}$ is countable, both $\sigma$-algebras coincide, but in general the cylinder $\sigma$-algebra is strictly smaller. We will use the cylinder $\sigma$-algebra to define universally measurable maps $f:\mathbb{R}^{\mathcal{P}}\rightarrow\mathbb{R}^{\mathcal{P}}$. The reason for this choice is the following characterization:
Proposition 4.4.
Let $X$ be a Hausdorff space and $Y=\prod_{i\in I}Y_{i}$ be any product of measurable spaces $(Y_{i},\Sigma_{i})$ for $i\in I$. Let $\Sigma_{Y}$ be the cylinder $\sigma$-algebra generated by the measurable spaces $(Y_{i},\Sigma_{i})$. The following are equivalent for $f:X\rightarrow Y$.:
@@
The preceding proposition says that a transition map is universally measurable if and only if it is universally measurable on all its features; in other words, we can check measurability of a transition just by checking measurability feature by feature. This is the same as in the Baire class-1 case (compare with Proposition 1.10).
The main result in section 3 is that, as long as we work with countably many features, PAC-learning (or NIP) corresponds to relative compactness in the space of Baire class-1 functions. The following result (which does not assume countability of the number of features) gives an analogous characterization of the NIP in terms of universal measurability:
Theorem 4.5 (Bourgain-Fremlin-Talagrand, Theorem 2F in [BFT78]).
@@
This result allows us to formalize the concept of a deep computation being Monte Carlo computable, which we define below (Definition 4.6). To motivate the definition, let us first recall two facts:
@@
    Littlewoood’s second principle states that every Lebesgue measurable function is “nearly continuous”. The formal statement of this, which is Luzin’s theorem, is that if $(X,\Sigma,\mu)$ a Radon measure space and $Y$ is a second-countable topological space (e.g., $Y=\mathbb{R}^{\mathcal{P}}$ with $\mathcal{P}$ countable ) equipped with a Borel algebra, then any given $f:X\to Y$ is measurable if and only if for every $E\in\Sigma$ and every $\varepsilon>0$ there exists a closed $F\subseteq E$ such that the restriction $f|_{F}$ is continuous and $\mu(E\backslash F)<\varepsilon$.
@@
    Computability of deep computations is characterized in terms of continuous extendibility of computations. This is at the core of [ADI+24].
@@
Let $(L,\mathcal{P},\Gamma)$ be a CCS. We say that a transition $f:L\rightarrow L$ is universally Monte Carlo computable if and only if there exists $\tilde{f}:\mathcal{L}_{\text{sh}}\rightarrow\mathcal{L}_{\text{sh}}$ extending $f$ such that for every sizer $r_{\bullet}$ there is a sizer $s_{\bullet}$ such that the restriction $\tilde{f}|_{\mathcal{L}[r_{\bullet}]}:\mathcal{L}[r_{\bullet}]\rightarrow\mathcal{L}[s_{\bullet}]$ is universally measurable, i.e., $\pi_{P}\circ\tilde{f}|_{\mathcal{L}[r_{\bullet}]}:\mathcal{L}[r_{\bullet}]\rightarrow[-s_{P},s_{P}]$ is $\mu$-measurable for every Radon probability measure $\mu$ on $\mathcal{L}[r_{\bullet}]$ and $P\in\mathcal{P}$.
Remark 4.7.
Condition (2) of Theorem  shows that to study measure-theoretic versions of NIP, we need only consider compact subsets of $X$. Now, every Radon measure on a compact space is finite; hence, by proper normalization, it can be treated as a probability measure. Therefore, in the context of Monte Carlo measurability, we focus on Radon probability measures rather than general Radon measures.
@@
Let $(L,\mathcal{P},\Gamma)$ be a CCS satisfying the Extendibility Axiom. Let $R$ be an exhaustive collection of sizers. Let $\Delta\subseteq\Gamma$ be $R$-confined. If $\pi_{P}\circ\Delta|_{L[r_{\bullet}]}$ satisfies the NIP for all $P\in\mathcal{P}$ and all $r_{\bullet}\in R$, then every deep computation in $\Delta$ is universally Monte Carlo computable.
@@
Fix $P\in\mathcal{P}$ and $r_{\bullet}\in R$. By the Extendibility Axiom, $\pi_{P}\circ\tilde{\Delta}|_{\mathcal{L}[r_{\bullet}]}$ is a set of pointwise bounded continuous functions on the compact set $\mathcal{L}[r_{\bullet}]$. Since $\pi_{P}\circ\tilde{\Delta}|_{L[r_{\bullet}]}=\pi_{P}\circ\Delta|_{L[r_{\bullet}]}$ has the NIP, so does $\pi_{P}\circ\tilde{\Delta}|_{\mathcal{L}[r_{\bullet}]}$ by Lemma 1.12. By Theorem 4.7, we have $\overline{\pi_{P}\circ\tilde{\Delta}|_{\mathcal{L}[r_{\bullet}]}}\subseteq M_{r}(\mathcal{L}[r_{\bullet}])$ for all $r_{\bullet}\in R$ and $P\in\mathcal{P}$. Let $f\in\overline{\Delta}$ be a deep computation. Write $f=\mathcal{U}\lim_{i}\gamma_{i}$ as an ultralimit of computations in $\Delta$. Define $\tilde{f}:=\mathcal{U}\lim_{i}\tilde{\gamma_{i}}$. Then, $\tilde{f}:\mathcal{L}_{\text{sh}}\rightarrow\mathcal{L}_{\text{sh}}$ extends $f$. Since $\Delta$ is $R$-confined we have that $f:L[r_{\bullet}]\rightarrow L[r_{\bullet}]$ and $\tilde{f}:\mathcal{L}[r_{\bullet}]\rightarrow\mathcal{L}[r_{\bullet}]$ for all $r_{\bullet}\in R$. Lastly, note that for all $r_{\bullet}\in R$ and $P\in\mathcal{P}$ we have that $\pi_{P}\circ\tilde{f}|_{\mathcal{L}[r_{\bullet}]}\in\overline{\pi_{P}\circ\tilde{\Delta}|_{\mathcal{L}[r_{\bullet}]}}\subseteq M_{r}(\mathcal{L}[r_{\bullet}])$. ∎
Question 4.9.
@@
4.2. Talagrand stability and Monte Carlo computability of deep computations
There is another notion closely related to NIP, introduced by Talagrand in [TAL84] while studying Pettis integration. Suppose that $X$ is a compact Hausdorff space and $A\subseteq\mathbb{R}^{X}$. Let $\mu$ be a Radon probability measure on $X$. Given a $\mu$-measurable set $E\subseteq X$, a positive integer $k$ and real numbers $a<b$. we write:
$D_{k}(A,E,a,b)=\bigcup_{f\in A}\{x\in E^{2k}:f(x_{2i})\leq a,\hskip 2.84526ptf(x_{2i+1})\geq b\hskip 2.84526pt{\text{ for all }i<k}\}.$
@@ 4.2. Talagrand stability and Monte Carlo computability of deep computations
where $\mu^{*}$ denotes the outer measure (we need to work with outer measure since the sets $D_{k}(A,E,a,b)$ need not be $\mu$-measurable). The inequality certainly holds when $A$ is a countable set of continuous (or $\mu$-measurable) functions.
The main result of this section is that accumulation points of standard computations (i.e., deep computations) are Monte Carlo computable; this is Theorem 4.14 below. We now prove that deep computations that are Talagrand $\mu$-stable are $\mu$-measurable. But first, let us state the following useful characterization of measurable functions (compare with Fact 1.2):
Fact 4.10 (Lemma 413G in [FRE03]).
Suppose that $(X,\Sigma,\mu)$ is a measure space and $\mathcal{K}\subseteq\Sigma$ is a collection of measurable sets satisfying the following conditions:
  1. (1)
    $(X,\Sigma,\mu)$ is complete, i.e., for all $E\in\Sigma$ with $\mu(E)=0$ and $F\subseteq E$ we have $F\in\Sigma$.
  2. (2)
    $(X,\Sigma,\mu)$ is semi-finite, i.e., for all $E\in\Sigma$ with $\mu(E)=\infty$ there exists $F\subseteq E$ such that $F\in\Sigma$ and $0<\mu(F)<\infty$.
  3. (3)
    $(X,\Sigma,\mu)$ is saturated, i.e., $E\in\Sigma$ if and only if $E\cap F\in\Sigma$ for all $F\in\Sigma$ with $\mu(F)<\infty$.
  4. (4)
    $(X,\Sigma,\mu)$ is inner regular with respect to $\mathcal{K}$, i.e., for all $E\in\Sigma$ μ(E)=sup{μ(K):$K$ and KE}.
(In particular, if $X$ is Hausdorff, $\mu$ is a Radon measure on $X$ and $\mathcal{K}$ is the collection of compact subsets of $X$, all these conditions hold). Then, $f:X\rightarrow\mathbb{R}$ is measurable if and only if for every $K\in\mathcal{K}$ with $0<\mu(K)<\infty$ and $a<b$, either $\mu^{*}(P)<\mu(K)$ or $\mu^{*}(Q)<\mu(K)$ where $P=\{x\in K:f(x)\leq a\}$ and $Q=\{x\in K:f(x)\geq b\}$.
The following technical lemma will be instrumental to prove of Proposition 4.13, which, in turn, will yield the main result of the subsection, Theorem 4.14.
@@
First, we claim that a subset of a $\mu$-stable set is $\mu$-stable. To see this, suppose that $A\subseteq B$ and $B$ is $\mu$-stable. Fix any $\mu$-measurable $E\subseteq X$ of positive measure and $a<b$. Let $k\geq 1$ such that
$(\mu^{2k})^{*}(D_{k}(B,E,a,b))<(\mu(E))^{2k}.$
Since $A\subseteq B$, we have $D_{k}(A,E,a,b)\subseteq D_{k}(B,E,a,b)$; therefore,
$(\mu^{2k})^{*}(D_{k}(A,E,a,b))\leq(\mu^{2k})^{*}(D_{k}(B,E,a,b))<(\mu(E))^{2k}.$
We now show that $\overline{A}$ is $\mu$-stable. Fix $E\subseteq X$ measurable with positive measure and $a<b$. Let $a^{\prime}<b^{\prime}$ be such that $a<a^{\prime}<b^{\prime}<b$. Since $A$ is $\mu$-stable, let $k\geq 1$ be such that
$(\mu^{2k})^{*}(D_{k}(A,E,a^{\prime},b^{\prime}))<(\mu(E))^{2k}.$
If $x\in D_{k}(\overline{A},E,a,b)$, then there is $f\in\overline{A}$ such that $f(x_{2i})\leq a<a^{\prime}$ and $f(x_{2i+1})\geq b>b^{\prime}$ for all $i<k$. By definition of pointwise convergence topology, there exists $g\in A$ such that $g(x_{2i})<a^{\prime}$ and $g(x_{2i+1})>b^{\prime}$ for all $i<k$. Hence, $x\in D_{k}(A,E,a^{\prime},b^{\prime})$. We have shown that $D_{k}(\overline{A},E,a,b)\subseteq D_{k}(A,E,a^{\prime},b^{\prime})$; hence,
$(\mu^{2k})^{*}(D_{k}(\overline{A},E,a,b))\leq(\mu^{2k})^{*}(D_{k}(A,E,a^{\prime},b^{\prime}))<(\mu(E))^{2k}.$
It suffices to show that $\overline{A}\subseteq\mathscr{M}^{0}(X,\mu)$. Suppose that there exists $f\in\overline{A}$ such that $f\notin\mathscr{M}^{0}(X,\mu)$. By fact 4.10, there exists a $\mu$-measurable set $E$ of positive measure and $a<b$ such that $\mu^{*}(P)=\mu^{*}(Q)=\mu(E)$ where $P=\{x\in E:f(x)\leq a\}$ and $Q=\{x\in E:f(x)\geq b\}$. Then, for any $k\geq 1$: $(P\times Q)^{k}\subseteq D_{k}(\{f\},E,a,b)$, so $(\mu^{2k})^{*}(D_{k}(\{f\},E,a,b))=(\mu^{*}(P)\mu^{*}(Q))^{k}=(\mu(E))^{2k}$. Thus, $\{f\}$ is not $\mu$-stable. However, we argued above that a subset of a $\mu$-stable set must be $\mu$-stable, so we have a contradiction. ∎
Definition 4.12.
We say that $A$ is universally Talagrand stable if $A$ is Talagrand $\mu$-stable for every Radon probability measure $\mu$ on $X$.
We first observe that universal Talagrand stability corresponds to a complexity class smaller than or equal to the NIP:
@@
By Theorem 4.7, it suffices to show that $A$ is relatively countably compact in $\mathscr{M}^{0}(X,\mu)$ for every Radon probability measure $\mu$ on $X$. Since $A$ is Talagrand $\mu$-stable for any such $\mu$, we have $\overline{A}\subseteq\mathscr{M}^{0}(X,\mu)$ by Lemma 4.11. In particular, AA is relatively countably compact in $\mathscr{M}^{0}(X,\mu)$. ∎
Corollary 4.14.
Let $(L,\mathcal{P},\Gamma)$ be a CCS satisfying the Extendibility Axiom. If $\pi_{P}\circ\Delta|_{L[r_{\bullet}]}$ is universally Talagrand stable for all $P\in\mathcal{P}$ and all sizers $r_{\bullet}$, then every deep computation is universally Monte Carlo computable.
Proof.
This is a direct consequence of Proposition 4.13 and Theorem 4.8. ∎
In the context of deep computations, we have identified two ways to obtain Monte Carlo computability, namely, NIP/PAC and Talagrand stability. It is natural to ask whether these two notions are equivalent. The following results show that, even in the simple case of countably many computations, this question is sensitive to the set-theoretic axioms. On the one hand, it is consistent (with respect to the standard ZFC axioms of set theory) that these two classes are the same:
Theorem 4.15 (Talagrand, Theorem 9-3-1(a) in [TAL84]).
@@
(The assumption that $[0,1]$ is not the union of $\mathfrak{c}$ closed measure zero sets is a consequence of, for example, the Continuum Hypothesis.)
On the other hand, by fixing a particular well-known measure, the Lebesgue measure, we see that the other case is also consistent:
Theorem 4.16 (Fremlin, Shelah, [FS93]).
@@
Notice that the preceding two results apply to sets of measurable functions, a class of functions larger than the class of continuous functions. It is possible that the continuity assumption, i.e., the Extendibility Axiom, settles the problem. This is a problem that remains open for future work.